OpenStudy (anonymous):

Please do not answer directly, I just need help working the problem. (2x - 4 / 10) = (2x / 10) solve for x. First step?

OpenStudy (campbell_st):

so is the question 1. \[\frac{2x - 4}{10} = \frac{2x}{10}\] or 2. \[2x - \frac{4}{10} = \frac{2x}{10}\]

OpenStudy (anonymous):

1.

OpenStudy (anonymous):

The problem as presented has no solution. Are you sure it's correct?

OpenStudy (campbell_st):

ok... the 1st step for these type of questions is to cross multiply to remove the denominators 10(2x - 4) = 10(2x) then distribute and then solve but the equation has no solution...

OpenStudy (anonymous):

Oops I apologize here is the correct equation 2x - 4 / 10 = 2x / 20 Again sorry about that

OpenStudy (anonymous):

Find the lowest common denominator and multiply both sides by that number.

OpenStudy (campbell_st):

ok... so still cross multiply 20(2x -4) = 10(2x)

OpenStudy (anonymous):

How do I multiply if there's a variable?

OpenStudy (campbell_st):

well you multiply each term inside the brackets by the term outside \[20 \times 2x - 20 \times 4 = 10 \times 2x\]

OpenStudy (anonymous):

40x - 80 = 20x?

Nnesha (nnesha):

yep.

OpenStudy (anonymous):

I got 0.4 = x. Is that correct?

Nnesha (nnesha):

replace x by .4 in the original equation if you get both equal sides then your answer 1000% right check!

OpenStudy (anonymous):

Yes it is correct thx everyone

Nnesha (nnesha):

is that point 4(.4) or just 4 ?

OpenStudy (anonymous):

.4

Nnesha (nnesha):

how did you get that ??

OpenStudy (anonymous):

It was a long process... Did I do something wrong? Or should I explain?

Nnesha (nnesha):

yes .4 is not a right answer explain plx if you show ur work ... that would be better :-)

OpenStudy (anonymous):

40x - 80 = 20x Starting from here i add 80 to both sides to get rid of 80 on one side 40x - 80 + 80 = 20x + 80 40x = 100x Divide both sides by 100? I think this might be where I'm doing something wrong...

Nnesha (nnesha):

nice first step is right add 80 both side then at right side you got 20x +80 now here these are not like terms ... 20 have x but 80 does not have any variable so you can't combine them

OpenStudy (anonymous):

So I have 40x and 20x + 80 Subtract 20x from both sides?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

20x = 80 divide both sides by 20?

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Ah, 4. Thanks for that! I'll apply this to some others I'm having problems with.

Nnesha (nnesha):

np :-)